You own a bond portfolio and decide to hedge using futures on 10 year Treasury notes. The futures contract has a price of $119.10 and a multiplier of $1,000. You identify a bond for delivery and use it to calculate the hedge ratio. The Treasury bond has a modified duration of 7.9.

If market rates rise by 23 basis points, by how much would the Treasury note futures be expected to change in value? Be sure to include the correct sign and enter your result rounded to the nearest integer.
Answer #1

To.find the change in value we will use the Basis Point Value (BPV)

For this,we need to find the money duration of the treasury notes. Money duration = modified duration X market value

Money Duration = 7.9*(119.10*1000) = $940,890

BPV = money duration * 0.0001

= $94.089

If the market rates rise by 23 basis points, value will change by

= -23*94.089 = -$2,164.047 or -$2,164 (when the interest rates rise, treasury notes fall in value, hence the negative sign)